OpenStudy (anonymous):

v(t)=2t²-7t+11 find the expression of v(2t) isit written as this? v(2t)=2(2t)²-7(2t)+11 is the answer for this 8t²-14t+11 ? or 8t²-14t+5 ?

OpenStudy (anonymous):

v(2t)=2(2t)^2-7(2t)+11 =2(4t^2)-14t+11 =8t^2-14t+11